# T-FAKE: Synthesizing Thermal Images for Facial Landmarking

### v1

@@ 1 Introduction
Functional thermal imaging is an acquisition modality with growing importance in many areas from life sciences over medicine to biometrics. Due to the close relation of body temperature with physiological states, thermal imaging of faces emerged as an important tool for screening infectious diseases [45, 40, 4] and for the analysis of neurocognitive and affective states [21]. As a consequence, there exists a growing need for thermal images with annotated facial landmarks. However, in contrast to RGB images, annotated data for thermal images of faces is still rarely available. For example, the largest thermal datasets with facial landmark annotations contain only tens to hundreds of individuals [29, 12, 43, 3], while the largest datasets for face detection for RGB images reached the mark of a million faces (of more than 500k individuals) with the MegaFace benchmark already in 2016 [24]. Moreover, large synthetic RGB datasets, like FAKE [52] appeared to be valuable for landmark detection training.
@@
Generally, the detection of facial landmarks is an important downstream task for facial analysis. These 2D coordinates of corresponding points across faces are part of methods, e.g., for alignment of faces to extract regions of interest [53], the animation of 3D avatars [35], motion magnification [14, 13] or deepfakes [46]. The most common convention for facial landmark annotations uses 68 points across the face and face boundary, see, e.g. [48]. Even though the 68-point convention remains challenging for thermal images [42, 37], the importance of dense landmarkers with more than 300 points has been demonstrated [52, 53, 35].
@@
Faces in thermal images share structural similarities with RGB images. However, temporally varying environmental factors such as temperature and rain droplets, and physiological factors such as blood perfusion changes and vascular details lead to much higher variability in the appearance of faces within individuals [20]. Sweating, flushing, or outdoor rain can result in time-dependent texture changes. Features that are very stable in RGB images such as the contrast of the nose and the eyes can become inverted when the nose or the face is cooled down. Due to the structural similarities, landmarkers trained on RGB datasets were also used under specific thermal imaging conditions [12]. However, they show high failure rates and produce imprecise landmark predictions on challenging thermal images. Up to now, there exist only a few sparse thermal landmark detection methods [29, 37, 43, 33, 32], and we are unaware of a dense thermal or a multimodal RGB+Thermal landmark system. In this paper, we aim to address this gap by introducing the T-FAKE dataset, the first large-scale synthetic thermal dataset featuring both sparse and dense landmarks for facial analysis. We propose to thermalize a synthetic RGB dataset using our novel RGB2Thermal loss function, which comprises three key components: i) a supervised data term that controls the generation of thermal faces based on a small subset of RGB-thermal pairs, ii) a Wasserstein distance term that aligns the patch distributions of the generated synthetic thermal images with those of real thermal images, and iii) a term incorporating prior information on clinical temperature statistics for different facial regions. To validate the quality of our T-FAKE dataset, we employ a probabilistic landmark prediction method by minimizing the negative log-likelihood function of the landmarks which are assumed to follow a Gaussian distribution as proposed in [53]. Moreover, by introducing a learnable adapter between different landmark conventions, we unify various methods and datasets for evaluation.
@@
We thoroughly evaluate the landmark prediction method trained on our T-FAKE dataset on the established RGB dataset 300W [48] and the challenging thermal dataset CHARLOTTE [3], which includes controlled modulation of environmental parameters such as pose and resolution. Our model demonstrates superior performance on the thermal landmark prediction benchmark compared to previous methods while maintaining performance on the RGB images comparable to state-of-the-art RGB methods. Fig. 1 illustrates the performance of our model for sparse and dense landmarking. In summary, we provide the following contributions:
@@
  • A novel RGB2Thermal loss to facilitate thermal image synthesis overcoming limited lab-recorded training data;
@@ 2 Related Work
ARL-MMFD [54] 2019 (request) 6 ?/111 250, 500, 750
DRIVE-IN [12] 2021 68 and 468 ?/436 $\sim$80-150
ARL-VTF [43] 2021 68 ?/395 210
@@
CHARLOTTE [3] 2022 73 (43 profile) 7684/10 100-660
T-FAKE (ours) 2024 70 and 468 200k/2k $\sim$80-150
@@
Facial RGB landmarking. The most widely used annotation convention defines 68-point landmarks which are available for many facial alignment datasets [27, 55, 6]. Sometimes points on the eye or other details are additionally added [52, 3]. While an in-depth review of facial RGB landmarking is beyond the scope of this work, we introduce those that are relevant to thermal landmarking methods. Kazemi et al. [23] introduced a traditional computer vision method for landmark prediction that is still widely used via the dlib library [25]. Kowalski et al. [30] proposed the deep alignment network (DAN) which uses convolutional neural networks and iteratively refines landmark predictions. Around the same time, Bulat et al. [6] proposed the facial alignment network (FA). Both methods predict heat maps of landmark positions from which coordinates are computed. Explicit integration of uncertainty quantification into a deep landmark prediction model via the LUVLi loss has been proposed by Kumar et al. [31]. In recent years, the quality of image annotation became a limiting factor of modern algorithms which led to the emergence of synthetic data with high-quality ground truth [39]. Wood et al. [52, 53] trained landmarkers on synthetic data that outperform traditional approaches and proposed the FAKE dataset with 100k frames of 2k individual faces.
@@
Thermal landmarking. Many tailored approaches for thermal landmarking use similar architectures as RGB landmarkers and are mainly developed alongside the publication of new datasets, see Table 1. Kuzdeuov et al. [32] released the TFW dataset with a total number of 9982 images from 147 subjects annotated with 5 facial landmarks and bounding boxes and trained YOLOv5 networks for landmark prediction. The SF-TL54 dataset [33] consists of 54 landmarks and 2556 frames of 142 subjects and has been used for landmark prediction with a dlib and a U-net model. Kopaczka et al. [29] also use a dlib model for their AACHEN dataset with annotated 2935 images of 90 subjects and 68-point landmarks. A highly structured dataset is the CHARLOTTE thermal face dataset with 10 subjects and 10k images by Ashrafi et al. [3], who annotated their dataset with 72-point landmarks. In contrast to these publicly available datasets, datasets with limited or no public availability include the the ARL-VTF dataset [43] (500k frames of 395 subjects) and a dataset recorded in SARS-CoV-2 drive-in stations (DRIVE-IN) [12] (six minutes recordings at $50$Hz of 436 subjects). The authors of the DRIVE-IN dataset found that a stack of pre-processing functions already leads to a good performance of existing RGB landmark models on thermal images. Similarly, Poster et al. [44] proposed to use transfer learning on a siamese DAN to improve thermal landmark prediction with RGB facial landmarks. Mallat et al. [37] used RGB2Thermal image translation (‘thermalization’), active appearance models, and deep alignment networks of RGB datasets for thermal landmark prediction. In contrast to the previous work, our dataset contains thermal images of 2k individuals that are annotated with sparse 70 as well as dense 478 point landmarks.
@@
Thermalization. In addition to thermal landmark datasets, various datasets of paired and aligned thermal and RGB images without annotations exist. These can be employed for the training of thermalization models. Mallat et al. thermalized facial images by minimizing a perceptual loss [37] using the VIS-TH dataset [36]. Similarly, generative adversarial networks (GANs) have been used for thermalization [54, 20] using the NCKU-VTF [20] dataset and the ARL-MMDF [54] dataset containing facial image data. In contrast to the limited public availability of these two datasets, a publicly available dataset with paired facial images is the multi-modal disguise Sejong face database (SEJONG) with 100 subjects, 15 disguise variations and 4849 RGB+Thermal image pairs by Cheema and Moon [7]. Thermalization of general scenes has been achieved by Kniaz et al. [26] with a modified Pix2Pix GAN, where they used the publicly available ThermalWorld dataset with 5098 image pairs displaying various object classes. In contrast to previous works, we employ a semi-supervised RGB2Thermal loss that incorporates a semantic segmentation to improve generalization.
@@ 3 T-FAKE Dataset
Recently, the success of landmarkers trained solely on the synthetic facial landmarking dataset FAKE [52] has demonstrated the potential of synthetic data for landmarking. We aim to transfer this success story to the largely overlooked thermal domain and the multimodal setting by introducing our synthetic thermal landmarking dataset T-FAKE featuring 2k unique subjects.
In particular, we train a semi-supervised image translation model for the thermalization of RGB images using real and synthetic data. To account for variations in the data in terms of poses, hairstyles, and expressions, we employ a semi-supervised loss.
@@ 3.1 Thermalization
To generate our large-scale synthetic T-FAKE dataset, we propose a thermalization framework. First, we train our network in a semi-supervised manner based on training pairs of RGB+Thermal images $(\mathbf{X}_{\text{ RGB}},\mathbf{X}_{\text{T}})=(X_{\text{ RGB}}^{n},X_{\text{T}}^{n})_{n=1}^{N}$ from the lab-recorded SEJONG dataset. Further, we use the synthetic RGB images from FAKE $\mathbf{X}_{\text{FAKE}}=(X_{\text{FAKE}}^{m})_{m=1}^{M}$ for which no thermal counterpart exits. Indeed the thermalization of these data appears to be difficult due to the large differences between real-world lab-recorded and synthetic in-the-wild data when facing thermalization tasks. For RGB2Thermal translation, we optimize a U-net [47] $T_{\theta}$ by minimizing the loss function
@@
This loss allows us to simultaneously train a domain-adaptive image translation model from paired and unpaired data. The first summand (1) penalizes the deviation of the generated thermal image from the true one and serves as a reconstruction loss for known image pairs. The second summand (2) exploits the Wasserstein distance on the patch distribution of thermalized FAKE images and natural thermal images. This style loss ensures that a generated image displays the thermal image statistics. To be more precise, recall that the squared Wasserstein-2 distance $\mathcal{W}_{2}^{2}$ between empirical measures $\mu=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$ and $\nu=\frac{1}{L}\sum_{l=1}^{L}\delta_{y_{l}}$ is given by
@@
$\mathcal{W}_{2,E}^{2}(\mu,\nu)$ $:=\inf_{\pi\in\Pi}\big{\{}F(\pi)+\lambda_{E}\sum_{k=1}^{K}\sum_{l=1}^{L}\pi_{k,l}\log\left(\pi_{k,l}\right)\big{\}}$
with some (small) regularization $\lambda_{E}>0$ has been proposed [9]. The new formulation and the resulting gradients can efficiently be calculated via the Sinkhorn algorithm [11, 41]. To use this loss for thermal translation, we focus on the patch distributions of the generated and real images. For that purpose, we now extract (overlapping) image patches $P_{i}X:=X[i_{1}:i_{1}+s,i_{2}:i_{2}+s]$, $i=(i_{1},i_{2})\in I$ of size $s\times s$ and consider the empirical measures generated by the patch distributions of $T_{\theta}(\mathbf{X}_{\text{FAKE}})$,
$\mu_{\text{FAKE}}:=\tfrac{1}{N}\tfrac{1}{|I|}\sum_{n=1}^{N}\sum_{i=1}^{|I|}\delta_{P_{i}\left(T_{\theta}(X_{\text{FAKE}}^{n})\right)},$
@@
During training, we apply a multiscale approach by applying this regularizer over multiple image scales and summing over the results. As a result, the patch statistics of downscaled images are incorporated into the empirical loss. The proposed patch distribution matching-based regularization was inspired by its application in related tasks such as image reconstruction [1, 41], style transfer [34], texture reconstruction [18, 2], and image translation [22].
@@
$\operatorname{R}\left(T_{\theta}(\mathbf{X}_{\text{FAKE}})\right)=\sum_{m=1}^{M}\sum_{i=1}^{18}\omega_{i}(\overline{S_{i}(T_{\theta}(X_{\text{FAKE}}^{m})})-T_{i})^{2},$
where the weights $\omega_{i}$ are proportional to the region sizes and $\overline{S_{i}}$ is the mean temperature in the $i$-th region. Note that facial temperatures depend on the environmental conditions. Hence, we train $T_{\theta}$ two times with two different reference temperature sets to simulate ‘cold’ and ‘warm’ environmental conditions. The different regularizers are visualized in Fig. 2 (left) and the segmentation masks are shown in Fig. 3 (bottom).
Predictions of the normalized temperature with our model compared to a model without regularization are shown in Fig. 3. The advantages of our regularization in a domain-adapted setting are clearly visible. The use of regularization reduces image artifacts. The patch regularizer (2) promotes natural texture synthesis and because of our segmentation regularizer (3) temperature differences between facial areas are more pronounced. Hence, our method allows for the adaptation of images with shadows, various skin colors, and distinctive facial expressions. In particular, such features may be underrepresented in facial thermal imaging datasets recorded in lab conditions. Note that the resulting domain adaptation needs for synthetic data and the access to ground truth segmentations differentiates our task from the thermalization tasks in [26, 54, 37, 20].
Figure 3: Frames from the FAKE dataset [52] before (first row) and after thermalization without any regularization (second row) and our domain-adaptive model (third row). The last row displays the segmentation masks with gray levels which are proportional to the reference temperature values used in regularizer (3). The first three images exemplify the ‘cold’ setup with colder noses and the last two the ‘warm’ setup with warmer noses.

### v2

@@ 1 Introduction
Functional thermal imaging is an acquisition modality with growing importance in many areas from life sciences over medicine to biometrics. Due to the close relation of body temperature with physiological states, thermal imaging of faces emerged as an important tool for screening infectious diseases [54, 46, 4] and for the analysis of neurocognitive and affective states [25]. As a consequence, there exists a growing need for thermal images with annotated facial landmarks. However, in contrast to RGB images, annotated data for thermal images of faces is still rarely available. For example, the largest thermal datasets with facial landmark annotations contain only tens to hundreds of individuals [34, 13, 52, 3], while the largest datasets for face detection for RGB images reached the mark of a million faces (of more than 500k individuals) with the MegaFace benchmark already in 2016 [29]. Moreover, large synthetic RGB datasets, like FAKE [61], appeared to be valuable for sparse and dense landmark detection training.
@@
Generally, the detection of facial landmarks is an important pre-processing task for further facial analysis. These 2D coordinates of corresponding points across faces are part of methods, e.g., for alignment of faces to extract regions of interest [62], the animation of 3D avatars [40], motion magnification [15, 14] or deepfakes [55]. The most common convention for facial landmark annotations uses 68 points across the face and face boundary, see, e.g. [58]. Even though the 68-point convention remains challenging for thermal images [51, 42], there is additional demand for dense landmarkers with more than 300 points. [61, 62, 40].
@@
Faces in thermal images share structural similarities with RGB images. However, temporally varying environmental factors such as temperature and rain droplets, and physiological factors such as blood perfusion changes and vascular details lead to much higher variability in the appearance of faces within individuals [22]. Sweating, flushing, or outdoor rain can result in time-dependent texture changes. Features that are very stable in RGB images such as the contrast of the nose and the eyes can become inverted when the nose or the face is cooled down. Due to the structural similarities, landmarkers trained on RGB datasets were also used under specific thermal imaging conditions [13]. However, they show high failure rates and produce imprecise landmark predictions on challenging thermal images. Up to now, there exist only a few sparse thermal landmark detection methods [34, 42, 52, 38, 37], and we are unaware of dense thermal or a multimodal RGB+Thermal landmarkers.
In this paper, we aim to address this gap by introducing the T-FAKE dataset, the first large-scale synthetic thermal dataset featuring both sparse and dense landmarks for facial analysis. We propose to thermalize a synthetic RGB dataset using our novel RGB2Thermal loss function, which comprises three key components: i) a supervised data term that controls the generation of thermal faces based on a small subset of RGB-thermal pairs, ii) a Wasserstein distance term that aligns the patch distributions of the generated synthetic thermal images with those of real thermal images, and iii) a term incorporating prior information on clinical temperature statistics for different facial regions. This loss enables our model to generalize beyond available lab-condition data to ‘in-the-wild’ images. Figure 4 shows T-FAKE samples and illustrates the advantage of our semi-supervised model over a supervised model trained solely with lab-condition images. To validate the quality of our T-FAKE dataset, we employ a probabilistic landmark prediction method by minimizing the negative log-likelihood function of the landmarks which are assumed to follow a Gaussian distribution as proposed in [62]. Moreover, by introducing a learnable adapter between different landmark conventions, we unify various methods and datasets for evaluation.
@@
We thoroughly evaluate the landmark prediction method trained on our T-FAKE dataset on the established RGB dataset 300W [58] and the challenging thermal dataset CHARLOTTE [3], which includes controlled modulation of temperature, pose, and resolution. Our model demonstrates superior performance on the thermal landmark prediction benchmark compared to previous methods while maintaining performance on the RGB images comparable to state-of-the-art RGB methods. Fig. 2 illustrates the performance of our model for sparse and dense landmarking. In summary, we provide the following contributions:
@@
  • A novel RGB2Thermal loss to facilitate thermal image synthesis overcoming lab-recorded training data with limited image variability;
@@ 2 Related Work
Facial RGB landmarking. Different annotation conventions exist for labeled landmarking datasets. An important distinction is between 2D annotations and 3DA-2D annotations which project 3D landmarks onto images [11]. 2D is closer to human perception, but 3DA-2D enables annotation of occluded face regions. The most widely used annotation convention defines 68-point 2D landmarks available for many facial alignment datasets [19, 32, 67, 6]. Sometimes points on the eye or other details are additionally added [61, 3]. While an in-depth review of facial RGB landmarking is beyond the scope of this work, we introduce those that are relevant to thermal landmarking methods. Kazemi et al. [28] introduced a popular traditional computer vision method for landmark prediction via the dlib library [30]. Kowalski et al. [35] proposed the deep alignment network (DAN) which uses convolutional neural networks and iteratively refines predictions. Around the same time, Bulat et al. [6] proposed the facial alignment network (FA). Both methods predict heat maps of landmark positions Explicit integration of uncertainty quantification into a landmark prediction model via the LUVLi loss has been proposed by Kumar et al. [36]. Recently, image annotation quality became a limiting factor which led to the emergence of synthetic data with high-quality ground truth [44]. Wood et al. [61, 62] trained 3DA-2D landmarkers with excellent performance using their proposed synthetic FAKE dataset with 100k frames of at least 2k individuals.
@@
ARL-MMFD [63] 2019 (request) 6 888/111 250, 500, 750
DRIVE-IN [13] 2021 68 and 478 $\sim$7-8M/436 $\sim$80-150
ARL-VTF [52] 2021 6 $500$k/395 210
@@
CHARLOTTE [3] 2022 72 (43 profile) 7684/10 100-660
T-FAKE (ours) 2024 70 and 478 200k/$\geq$2k $\sim$80-150
@@
Thermal landmarking. Many tailored approaches for thermal landmarking use similar architectures as RGB landmarkers and are mainly developed alongside the publication of new datasets, see Table 1. Kuzdeuov et al. [37] released the TFW dataset with with 5 facial landmarks and bounding boxes and trained YOLOv5 networks for landmark prediction. The SF-TL54 dataset [38] has been used for 2D landmark prediction with a dlib and a U-net model. Kopaczka et al. [34] also use a dlib model for their AACHEN dataset with 68-point 2D landmarks. A highly structured 2D landmarking dataset is the CHARLOTTE thermal face dataset with controlled temperature changes by Ashrafi et al [3] It contains around 10k images in total and 7684 images with annotations. In contrast to these publicly available datasets, datasets with limited or no public availability include the the ARL-VTF dataset [52] and a dataset recorded in SARS-CoV-2 drive-in stations (DRIVE-IN) [13] (six minutes recordings at $50$Hz of 436 subjects). The authors of the DRIVE-IN dataset found that a stack of pre-processing functions already leads to a good performance of existing RGB landmark models on thermal images. Similarly, Poster et al. [53] proposed to use transfer learning on a siamese DAN to improve thermal landmark prediction with RGB facial landmarks. Mallat et al. [42] used RGB2Thermal image translation (‘thermalization’), active appearance models, and deep alignment networks of RGB datasets for thermal landmark prediction. In contrast to the previous work, our dataset contains thermal images of at least 2k individuals that are annotated with sparse 70 3DA-2D as well as dense 478 2D point landmarks.
@@
Thermalization. In addition to thermal landmark datasets, various datasets of paired and aligned thermal and RGB images without annotations exist. These can be employed for the training of thermalization models. In contrast to the rich literature on Thermal2RGB synthesis [56, 64, 45], the literature on RGB2Thermal image synthesis using such datasets is more limited. Mallat et al. thermalized facial images by minimizing a perceptual loss [42] using the VIS-TH dataset [41]. Similarly, generative adversarial networks (GANs) have been used for Thermal2RGB synthesis [63, 22] using the NCKU-VTF [22] dataset and the ARL-MMDF [63] dataset containing facial image data. In contrast to the limited public availability of these two datasets, a publicly available dataset with paired facial images is the multi-modal disguise Sejong face database (SEJONG) with 100 subjects, 15 disguise variations and 4849 RGB+Thermal image pairs by Cheema and Moon [7]. Note that all these datasets and resulting models are mostly restricted to frontal views of faces with controlled lighting. Thermalization of general scenes has been achieved by Kniaz et al. [31] with a modified Pix2Pix GAN, where they used the publicly available ThermalWorld dataset with 5098 triplets of RGB images, thermal images, and segmentation masks. In contrast to previous works, we employ a semi-supervised RGB2Thermal loss using a semantic segmentation to improve generalization beyond lab-condition frontal views to ‘in-the-wild’ images.
@@ 3 T-FAKE Dataset
Recently, the success of facial landmarkers trained solely on the FAKE dataset [61], synthetic facial landmarking dataset, has demonstrated the potential of synthetic data for the task. We aim to transfer this success story to the largely overlooked thermal domain and the multimodal setting by introducing our synthetic thermal landmarking dataset T-FAKE featuring at least 2k unique subjects. In particular, we train a semi-supervised image translation model for the thermalization of RGB images using real and synthetic data. To account for data variations in terms of poses, shadows, and expressions, we employ a semi-supervised loss.
@@ 3.1 Thermalization
To generate our large-scale synthetic T-FAKE dataset, we propose a thermalization framework. First, we train our network in a semi-supervised manner based on training pairs of RGB+Thermal images $(\mathbf{X}_{\text{ RGB}},\mathbf{X}_{\text{T}})=(X_{\text{ RGB}}^{n},X_{\text{T}}^{n})_{n=1}^{N}$ from the lab-recorded SEJONG dataset. Further, we use the synthetic RGB images from FAKE $\mathbf{X}_{\text{FAKE}}=(X_{\text{FAKE}}^{m})_{m=1}^{M}$ for which no thermal counterpart exits. The thermalization of these data appears to be difficult due to the large differences between real-world lab-recorded and synthetic in-the-wild data when facing thermalization tasks. For RGB2Thermal translation, we optimize a U-net [57] $T_{\theta}$ by minimizing the loss function
@@
This loss allows us to simultaneously train a domain-adaptive image translation model from paired and unpaired data. The first summand (1) penalizes the deviation of the generated thermal image from the true one and serves as a reconstruction loss for known image pairs. The second summand (2) exploits the Wasserstein distance on the patch distribution of thermalized FAKE images and natural thermal images. This style loss ensures that a generated image displays the thermal image statistics. To be more precise, recall that the squared Wasserstein-2 distance $\mathcal{W}_{2}^{2}$ between empirical measures $\mu=\frac{1}{K}\sum_{k=1}^{K}\delta_{x_{k}}$ and $\nu=\frac{1}{L}\sum_{l=1}^{L}\delta_{y_{l}}$ for Dirac probability measures $\delta_{\bullet}$ is given by
@@
$\mathcal{W}_{2,E}^{2}(\mu,\nu)$ $:=\inf_{\pi\in\Pi}\big{\{}F(\pi)+\lambda_{E}\sum_{k=1}^{K}\sum_{l=1}^{L}\pi_{k,l}\log\left(\pi_{k,l}\right)\big{\}}$ (5)
with some (small) regularization $\lambda_{E}>0$ has been proposed [9]. This term and its gradients can efficiently be calculated via the Sinkhorn algorithm [12, 50]. To use this loss for thermal translation, we focus on the patch distributions of the generated and real images. For that purpose, we now extract (overlapping) image patches by defining the patch extraction operator $P_{i}X:=X[i_{1}:i_{1}+s,i_{2}:i_{2}+s]$, $i=(i_{1},i_{2})\in I$ of size $s\times s$ and considering the empirical measures generated by the patch distributions of $T_{\theta}(\mathbf{X}_{\text{FAKE}})$,
$\mu_{\text{FAKE}}:=\tfrac{1}{{\color[rgb]{0,0,0}M}}\tfrac{1}{|I|}\sum_{{\color[rgb]{0,0,0}m}=1}^{{\color[rgb]{0,0,0}M}}\sum_{i=1}^{|I|}\delta_{P_{i}\left(T_{\theta}(X_{\text{FAKE}}^{n})\right)},$ (6)
@@
During training, we apply a multiscale approach by applying this regularizer over multiple image scales and summing over the results (see Table 4). As a result, the patch statistics of downscaled images are incorporated into the empirical loss. The proposed patch distribution matching-based regularization was inspired by its application in related tasks such as image reconstruction [1, 50], style transfer [39], texture reconstruction [20, 2], and image translation [26].
@@
$\operatorname{R}\left(T_{\theta}(\mathbf{X}_{\text{FAKE}})\right)=\sum_{m=1}^{M}\sum_{i=1}^{18}\omega_{i}(\overline{S_{i}(T_{\theta}(X_{\text{FAKE}}^{m})})-T_{i})^{2},$ (8)
where the weights $\omega_{i}$ are proportional to the region sizes and $\overline{S_{i}}$ is the mean temperature in the $i$-th region. Note that facial temperatures depend on the environmental conditions. Hence, we train $T_{\theta}$ two times with two different reference temperature sets to simulate ‘cold’ and ‘warm’ environmental conditions. The different regularizers are shown in Fig. 3 (left) and the segmentations in Fig. 4 (bottom).
Predictions of the normalized temperature with our model compared to a model without regularization are shown in Fig. 4. The advantages of our regularization in a domain-adapted setting are clearly visible. The use of regularization reduces image artifacts. The patch regularizer (2) promotes natural texture synthesis and because of our segmentation regularizer (3) temperature differences between facial areas are more pronounced. Hence, our method allows for the adaptation of images with shadows, various skin colors, and distinctive facial expressions. In particular, such features may be underrepresented in facial thermal imaging datasets recorded in lab conditions. Note that the resulting domain adaptation needs for synthetic data solved by a semi-supervised approach based on the access to ground truth segmentations differentiates our task from the RGB2Thermal tasks in [31, 42] and (supervised) Thermal2RGB tasks in [64, 45, 22]. The generalization limitations of supervised models are further discussed in Section 5.5, see Table 4, and the supplementary.